OpenStudy (anonymous):

what is z when z-2/9=1/9 what does z equal

OpenStudy (ajprincess):

Is ur question \[z-\frac{ 2 }{ 9 }=\frac{ 1 }{ 9}?\]

OpenStudy (ajprincess):

@jordanqweqwe

OpenStudy (aravindg):

if thats the case z=1/9+2/9=3/9=1/3